# Think-on-Graph 2.0: Deep and Interpretable Large Language Model Reasoning with Knowledge Graph-guided Retrieval

### v1

@@ Abstract.
XXX
journalyear: 2023booktitle: price: 15.00doi: 10.1145/3539618.3591766isbn: 978-1-4503-9408-6/23/07ccs: Information systems Users and interactive retrieval
1. Introduction
Figure 1. Comparison of traditional RAG (a), KG-based generation (b) and Graph-guided RAG (c). The example illustrates the limitations of pure semantic retrieval and pure knowledge graph-augmented frameworks in complex QA tasks, and the advantages of our proposed KG+RAG framework. a: The dataset size is too large in the semantic retrieval-based RAG paradigm, resulting in low information density. Also, traditional retrieval systems struggle to capture deep connections between facts, thus failing to focus on key points in the question. b: In the KG path inference-based LLM augmentation paradigm, the information provided by triples lacks both depth and details. Even information may be missing, due to the incompleteness of the KG. c: The proposed ToG2.0 combines the advantages of both approaches. The KG helps to understand deep connections between different facts and precisely narrows down the search scope, while entity context-based retrieval supplements the information missing in the knowledge graph’s triple path reasoning.
@@ 1. Introduction
Figure 2. The detailed structure of ToG2.0.
@@
2. Related Works
2.1. Retrieval Augmented Generation with Knowledge Graph
@@
KGs offer dynamic, explicit, and structured representations of knowledge. This structured knowledge representation is particularly beneficial for LLMs because it introduces a level of interpretability and precision in the knowledge that LLMs can access. Early approaches (Sun et al. (2020), Peters et al. (2019), Huang et al. (2024), Liu et al. (2020)) focused on embedding knowledge from KGs directly into the neural networks of LLMs. This embedding could occur either during the pretraining phase or the fine-tuning process. The intent was to infuse the models with rich, structured knowledge right from the foundational stages of model training. Despite the initial promise, integrating KGs directly into LLMs introduced challenges. As noted by Hu et al. (2023), this integration leads to a reduction in the natural explainability of the models. Additionally, it makes updating the knowledge base more complex and less efficient, as any change in the KG requires retraining or significant adjustments to the model. More recent studies(Jiang et al., 2024) have shifted towards using KGs to augment LLMs externally rather than embedding the knowledge directly into the models. Those approaches involve translating relevant structured knowledge from KGs into textual prompts that are then fed to LLMs. The process typically follows a fixed pipeline where extra information from KGs is retrieved to enhance the LLM prompts. The integration of Knowledge Graphs (KGs) with Large Language Models (LLMs) offers numerous advantages but also meets several distinct challenges and limitations, such as incompleteness and ambiguity as discussed in the last section.
@@ 3. Methodology
In complex reasoning, the implicit correlation between the question and effective intermediate clue sentences often goes unrecognized by both sparse retrieval models and dense pre-trained retrieval models. To solve this limitation, we prompt the LLM to formulate clue queries $q_{j}^{0}$ based on the current context for every topic entity $e_{j}$, which orients the next-step direction of exploring the relations and contexts. Using the question about Tencent mentioned above as the example again, based on the entity ”National People’s Congress?”, the LLM may generate a clue-query that suggests gathering information about their political roles or affiliations.
3.2. Reasoning with Graph-driven Knowledge Retrieval
@@
Our Model LLM WebQSP HotpotQA QALD-10-en FEVER
Tog2.0 GPT-3.5-turbo 81.13 40.91 54.05 58.54
Baseline
Vanilla GPT-3.5-turbo 74.55 28.89 42.04 52.10
CoT GPT-3.5-turbo 59.93 34.40 42.94 57.80
CoK GPT-3.5-turbo - 35.40 - 63.50
ToG GPT-3.5-turbo 76.20 26.30 50.20 52.70
Table 1. Performance of our method versus various baselines. The evaluation metric for FEVER is Accuracy, while the metric for WebQSP, HotpotQA, and QALD-10-en is Exact Match (EM).
@@
Tog2.0 (w/o TP, RC, clue_query) GPT-3.5-turbo 78.70 39.29 51.05 56.30
Tog2.0 (w/o TP, RC, clue_query) Llama-2-13b 76.22 29.15 48.64 49.17
Tog2.0 (w/o TP, clue_query) GPT-3.5-turbo 76.43 38.64 49.85 56.04
@@
Table 2. Ablation study: the influence of each module in Tog2.0 on the final performance. The evaluation metric for FEVER is Accuracy, while WebQSP & HotpotQA & QALD-10-en is Exact Match (EM). clue-query: query re-formulate. TP: topic prune. RC: relation prune combination.
@@
Examine and reasoning (ER): Following RP and EP, we give LLM carefully aggregated references, including $q$, $Q_{c}^{i}$, $\mathbf{P}^{i+1}$ and the top $L$ ($L\leq K$) chunks. Then the LLM is prompted to examine the logical coherence and the completeness of factual evidence. If the LLM believes it can answer the question, the iteration ends. If not, based on the question and the collected contextual clues, a new clue-query needs to be generated for the next round.
4. Experiments
4.1. Datasets and Metrics
@@ 4.5. Ablation Study
Additionally, clue-query also brought relatively consistent improvements across each dataset, indicating that adaptive query optimization can help the LLM better understand the tasks. We also tested the vanilla RAG process and the basic Tog2.0 process on Llama-2-13B. It can be seen that on a less capable LLM, Tog2.0 can bring a greater performance improvement. This suggests that Tog2.0 might be more adaptable. Weaker LLMs often encounter bottlenecks when handling complex tasks, while Tog2.0 utilizes knowledge graphs as clues to optimize the reasoning path and reduce task complexity. It then uses entity context to further guide the model to focus on relevant information, thereby improving task understanding and response accuracy. In contrast, GPT-3.5, due to its higher inherent capabilities, may not exhibit as significant a performance improvement because it is already close to its performance ceiling.

### v2

@@ Abstract
1 Introduction
Figure 1: Comparison of traditional RAG (a), KG-based generation (b), and Graph-guided RAG (c). The example illustrates the limitations of pure semantic retrieval and pure knowledge graph-augmented frameworks in complex QA tasks, and the advantages of our proposed KG+RAG framework. a: The dataset size is too large in the semantic retrieval-based RAG paradigm, resulting in low information density. Also, traditional retrieval systems struggle to capture deep connections between facts, thus failing to focus on key points in the question. b: In the KG path inference-based LLM augmentation paradigm, the information provided by triples lacks both depth and details. Even information may be missing, due to the incompleteness of the KG. c: The proposed ToG2.0 combines the advantages of both approaches. The KG helps to understand deep connections between different facts and precisely narrows down the search scope, while entity context-based retrieval supplements the information missing in the knowledge graph’s triple path reasoning.
@@ 1 Introduction
Figure 2: The detailed structure of ToG2.0.
@@
2 Related Works
2.1 Retrieval Augmented Generation with Knowledge Graph
@@
KGs offer dynamic, explicit, and structured representations of knowledge. This structured knowledge representation is particularly beneficial for LLMs because it introduces a level of interpretability and precision in the knowledge that LLMs can access. Early approaches (Sun et al., 2020; Peters et al., 2019; Huang et al., 2024; Liu et al., 2020) focused on embedding knowledge from KGs directly into the neural networks of LLMs. This embedding could occur either during the pretraining phase or the fine-tuning process. The intent was to infuse the models with rich, structured knowledge right from the foundational stages of model training. Despite the initial promise, integrating KGs directly into LLMs introduced challenges. As noted by Hu et al. (2023), this integration leads to a reduction in the natural explainability of the models. Additionally, it makes updating the knowledge base more complex and less efficient, as any change in the KG requires retraining or significant adjustments to the model. More recent studies (Jiang et al., 2024) have shifted towards using KGs to augment LLMs externally rather than embedding the knowledge directly into the models. Those approaches involve translating relevant structured knowledge from KGs into textual prompts that are then fed to LLMs. The process typically follows a fixed pipeline where extra information from KGs is retrieved to enhance the LLM prompts. The integration of Knowledge Graphs (KGs) with Large Language Models (LLMs) offers numerous advantages but also meets several distinct challenges and limitations, such as incompleteness and ambiguity as discussed in the last section.
@@ 3.1 Initializtion
In complex reasoning, the implicit correlation between the question and effective intermediate clue sentences often goes unrecognized by both sparse retrieval models and dense pre-trained retrieval models. To solve this limitation, we prompt the LLM to formulate clue queries $q_{j}^{0}$ based on the current context for every topic entity $e_{j}$, which orients the next-step direction of exploring the relations and contexts. Using the question about Tencent mentioned above as the example again, based on the entity ”National People’s Congress?”, the LLM may generate a supplementary query that suggests gathering information about their political roles or affiliations.
3.2 Reasoning with Graph-driven Knowledge Retrieval
@@
Table 1: Performance of our method versus various baselines. The evaluation metric for FEVER is Accuracy, while the metric for WebQSP, HotpotQA, and QALD-10-en is Exact Match (EM).
Model Tog2.0 Vanilla CoT CoK ToG
LLM GPT-3.5-turbo GPT-3.5-turbo GPT-3.5-turbo GPT-3.5-turbo GPT-3.5-turbo
WebQSP 81.13 74.55 59.93 - 76.20
HotpotQA 40.91 28.89 34.40 35.40 26.30
QALD-10-en 54.05 42.04 42.94 - 50.20
FEVER 58.54 52.10 57.80 63.50 52.70
Creak 93.5 89.7 90.1 - 93.8
Zero-Shot RE 91.0 27.7 28.8 - 88.0
Table 2: Ablation study: the influence of each module in Tog2.0 on the final performance. The evaluation metric for FEVER is Accuracy, while WebQSP & HotpotQA & QALD-10-en is Exact Match (EM). SQ: supplementary query. TP: topic prune. RC: relation prune combination.
@@
Tog2.0 (w/o TP&RC&SQ) GPT-3.5-turbo 78.70 39.29 51.05 56.30
Tog2.0 (w/o TP&RC&SQ) Llama-2-13b 76.22 29.15 48.64 49.17
Tog2.0 (w/o TP&SQ) GPT-3.5-turbo 76.43 38.64 49.85 56.04
@@
Examine and reasoning (ER): Following RP and EP, we give LLM carefully aggregated references, including $q$, $Q_{c}^{i}$, $\mathbf{P}^{i+1}$ and the top $L$ ($L\leq K$) chunks. Then the LLM is prompted to examine the logical coherence and the completeness of factual evidence. If the LLM believes it can answer the question, the iteration ends. If not, based on the question and the collected contextual clues, a new supplementary query needs to be generated for the next round.
4 Experiments
4.1 Datasets and Metrics
@@ 4.5 Ablation Study
Additionally, the supplementary query also brought relatively consistent improvements across each dataset, indicating that adaptive query optimization can help the LLM better understand the tasks. We also tested the vanilla RAG process and the basic Tog2.0 process on Llama-2-13B. It can be seen that on a less capable LLM, Tog2.0 can bring a greater performance improvement. This suggests that Tog2.0 might be more adaptable. Weaker LLMs often encounter bottlenecks when handling complex tasks, while Tog2.0 utilizes knowledge graphs as clues to optimize the reasoning path and reduce task complexity. It then uses entity context to further guide the model to focus on relevant information, thereby improving task understanding and response accuracy. In contrast, GPT-3.5, due to its higher inherent capabilities, may not exhibit as significant a performance improvement because it is already close to its performance ceiling.